OpenStudy (anonymous):

how do i solve 16/x-1=x-1/4? Its a proportion and i need to be walked through it

OpenStudy (saifoo.khan):

\[\frac{16}{x-1}=\frac{x-1}{4}\] Like that?

OpenStudy (anonymous):

\[\frac{16}{x-1}=\frac{x-1}{4}\] like that?

OpenStudy (anonymous):

lolololol!

hero (hero):

WoW

OpenStudy (saifoo.khan):

HAHA!

OpenStudy (anonymous):

saifoooooooooooooooooo! you sound like me now

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Copy Cat.

OpenStudy (anonymous):

or maybe (heaven forbid) i sound like you...

hero (hero):

You two are just weird. That's all

OpenStudy (saifoo.khan):

Im following your track @sir Satellite! ;)

OpenStudy (anonymous):

cross multiply

OpenStudy (saifoo.khan):

HAHA, Hero's Jealous!

OpenStudy (anonymous):

\[16\times 4=(x-1)(x-1)\]

hero (hero):

Saif you always assume I am.

OpenStudy (anonymous):

okay, did and ended up with 66= x^2 - 2x

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

so \[(x-1)^2=64\] \[x-1=8\] or \[x-1=-8\]

OpenStudy (saifoo.khan):

That's the fact @hero, u cant hide that.

OpenStudy (anonymous):

and therefore \[x=9\] or \[x=-7\]

OpenStudy (anonymous):

@monica don't multiply out

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you already have a perfect square so leave it at that

OpenStudy (anonymous):

if it was not a perfect square you would have to multiply out and solve a quadratic, but not in this case

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (saifoo.khan):

please hit good answer for Satellite.

OpenStudy (anonymous):

saifoooooo passes me third week of september (prediction)

OpenStudy (saifoo.khan):

LOL^

OpenStudy (saifoo.khan):

Lets C.. ;)